OpenStudy (anonymous):

Solve: x + 1 = −x + 8

OpenStudy (ghazi):

2x=7 x=7/2

OpenStudy (anonymous):

We want the variables, x , on one side and the constants on the other. x + 1 + x = -x + 8 +x Adding x to both sides 2x +1 = 8 Now we subtract 1 from both sides. 2x = 7 Divide by 2 x = 7/2